OpenStudy (anonymous):

How to solve an equation using factoring and quadratic formula ??? x^-19x+48 = 0 ???

OpenStudy (anonymous):

factoring (x-16)(x-3) = 0 so x = 16 or 3

OpenStudy (anonymous):

(x -16)(x-3) =0 x + 3 or 16

OpenStudy (anonymous):

You know how to solve using the Quadratic formula right?

OpenStudy (anonymous):

noo

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

The formula is \[x = (-b \pm \sqrt{b ^{2}-4ac}) / 2a\]

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

here b = -19 a = 1 c = 48 Solve :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

welcome jen